A dart player throws a dart horizontally at the wall from a distance of 8 meters. On its way to the target, the dart drops 0.32 meters. At what horizontal speed was the dart thrown?
Answer by Alan3354(69443)

A dart player throws a dart horizontally at the wall from a distance of 8 meters. On its way to the target, the dart drops 0.32 meters. At what horizontal speed was the dart thrown?
----------------
Using 9.8 m/sec/sec for gravity:
0.32 = 4.9t^2 where t = flight time
t =~ 0.2555 seconds
---------------
d = r*t
8 = r*t
r = 8/t =~ 31.3 meter/sec